OpenStudy (anonymous):

Which equation is a quadratic equation? Answer y + 3 = 5x + 10 y + 4x2 = (4x - 2)(x + 3) y - 8x = (x2 + 1)(9x - 7) y = 2x + 7x(x + 1)

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

it's not #3

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

y - 8x = (x^2 + 1)(9x - 7) y - 8x = 9x^3 - 7x^2 + 9x - 7 which is a cubic (and not a quadratic)

OpenStudy (apoorvk):

the equation with degree '2', when written in implicit form.

OpenStudy (anonymous):

ohhh... so its #2

jimthompson5910 (jim_thompson5910):

close, but again no

jimthompson5910 (jim_thompson5910):

y + 4x^2 = (4x - 2)(x + 3) y + 4x^2 = 4x^2 + 12x - 2x - 6 y = 4x^2 + 12x - 2x - 6 - 4x^2 y = 10x - 6 ... which is linear

OpenStudy (anonymous):

confused... then it has to be the last one... cause im pretty sure the first one isnt a quadratic

jimthompson5910 (jim_thompson5910):

you got it, you can either use the process of elimination, or you can do... y = 2x + 7x(x + 1) y = 2x + 7x^2 + 7x y = 7x^2 + 9x which is definitely a quadratic